# Tony Lindeberg

### v2

@@ 1 Introduction
For simplicity, we have here discarded a possible complementary variability with respect to an added translation vector $b$ in the image domain of the form $x^{\prime}={\cal A}\,x+b$, by without essential loss of generality assuming that all the multi-view observations of the same surface patch are focused on the same viewing point $P$ on the 3-D object, and with the origins of the coordinate systems in the image domain also corresponding to the projections of that same fixation point.
@@
A conceptual question that one could then ask is if the ability of perceptual system to derive seemingly stable representations of external objects in the 3-D environment would lead to constraints regarding the variabilities in the shapes of the receptive fields at the earliest levels in the visual processing hierarchy. In a theoretically motivated and axiomatically determined normative theory of visual receptive fields (Lindeberg 2013a; 2021), idealized models for the receptive fields of simple cells in the primary visual cortex (V1) have specifically been derived, that lead to provable covariance properties under spatial affine transformations, and regarding extensions to time-dependent spatio-temporal image data also provable covariance properties under Galilean transformations, to handle the variabilities in the relative motion between the objects in the world and the observer, see Lindeberg (2023; 2024c) for more in-depth treatments regarding the importance of covariance properties for visual receptive fields under geometric image transformations.
From a computational viewpoint, if we regard the earliest layers in the visual perception system as a computational engine, that infers successively more complex cues about the structure of the environment from the image measurements that reach the retina, then covariance properties of visual receptive fields do from a theoretical perspective constitute a highly useful notion. Specifically, if we design a computational algorithm for computing the local surface orientation of a smooth surface patch from binocular cues, then compared to not basing the computations on covariant spatial receptive fields, the use of provably affine-covariant visual receptive fields can improve the accuracy by an order of magnitude in estimates of the local surface orientation (see Tables 1–4 in Lindeberg and Gårding (1997)), by eliminating a source of error corresponding to a mismatch between the backprojected receptive fields to the tangent plane of the surface, from the two different perspective views, see Figure 2 for an illustration.
@@
  • first theoretically describing the different degrees of freedom in spatial affine transformations, in terms of a canonical parameterization, based on a closed-form factorization of the affine transformation matrix ${\cal A}$ very closely related to a singular value decomposition,
  • relating these degrees of freedom to the degrees of freedom in the parameters of the idealized receptive models of simple cells in the primary visual cortex, in terms of the generalized Gaussian derivative theory for visual receptive fields, and finally
  • relating these results to existing evidence regarding variabilities in the shapes of biological receptive fields established from neurophysiological measurements.
Specifically, we will after an initial overview of related work in Section 2, in Section 3 describe a decomposition of 2-D spatial affine image transformations, based on a closed-form decomposition of the affine transformation matrix in a way very closely related to a singular value decomposition, and which reveals the degrees of freedom of 2-D spatial affine image transformations in a very geometric manner.
@@
Based on this theoretical basis, we will then in Section 6 relate the degrees of freedom in the affine Gaussian derivative model to the degrees of freedom revealed by our proposed decomposition of 2-D affine transformations, in terms of explicit mappings between the parameters in our decomposition of 2-D spatial affine transformations and the parameters in the affine Gaussian derivative model. These connections will then be used in Section 7 for addressing the question of whether we can regard the receptive fields in the primary visual cortex as being able to span the degrees of freedom of 2-D affine transformations.
@@ 2 Relations to previous work
The receptive field shapes of simple cells have, in turn, been modelled mathematically in terms of, most commonly, either Gabor filters by Marcelja (1980), Jones and Palmer (1987a; 1987b), Porat and Zeevi (1988), Ringach (2002; 2004), or Gaussian derivatives by Koenderink and van Doorn (1984; 1987; 1992) Young and his co-workers (1987; 2001; 2001) and Lindeberg (2013a; 2021).
@@
The subject of this paper is to address this issue with regard to special case of spatial affine image transformations, and specifically based on a canonical decomposition of such 2-D spatial affine transformations, on a product form of a uniform scaling transformation, pure rotations, a non-uniform scaling transformation, and a normalization of the non-uniform scaling transformation to a preferred symmetry orientation in the image domain.
An earlier study of parameterizations of affine transformations, which we will build upon in this work, was presented in (Lindeberg 1995). In this paper, we extend that parameterization by first of all providing an explicit derivation in Appendix A.1, which was not presented in the original paper because of space limitations, and then analysing the properties of this decomposition for four basic classes of primitive affine transformations in Appendix A.2. We do additionally give a more general treatment regarding special cases for the parameters in the affine transformation model, which thereby increases the domain of applicability of the proposed parameterization to more general families of affine transformation matrices, that are reasonably near the unit matrix multiplied by a scalar scaling factor.
@@ 3 The degrees of freedom in 2-D spatial affine transformations
and parameterize the image coordinates as $x=(x_{1},x_{2})^{T}$ and $x^{\prime}=(x^{\prime}_{1},x^{\prime}_{2})^{T}$, where we are here specifically interested in the special case when the affine transformation matrix ${\cal A}$ is reasonably close to the unit matrix ${\cal I}$ multiplied by some scalar scaling factor.
@@ 3.1 Geometric interpretations of affine image deformations
As previously mentioned, when surface patterns of possibly moving smooth objects in the world are projected to the image plane, by either (i) the perspective mapping to a monocular image, (ii) the projective mapping between two binocular views, or (iii) the temporally integrated optic flow fields obtained by observing moving objects between at adjacent time moments, then the resulting image transformations or deformation fields can to first order of approximation be modelled as local affine transformations.
For this purpose, the decomposition in (LABEL:eq-aff-decomp) constitutes a geometrically very natural way to parameterize the degrees of freedom in such local affine transformations of the perspective/projective mappings, and which thus determine the variabilities of the image structures, that a visual system is exposed to, when observing sufficiently smooth surface patterns in the environment from different viewing directions relative to the objects in the environment.
@@ 3.1.1 Monocular locally linearized perspective projection
To make such relations more explicit, consider, for example, the monocular perspective projection of a smooth local surface patch onto a spherical camera. Then, in two local coordinate systems, with
@@
  • the second local coordinate system being in the tangent plane of the spherical camera, with the second local coordinate direction aligned with the backprojected tilt direction onto the resulting image plane parallel to the tangent plane of the surface,
the first-order locally linearized mapping from the tangent plane of the surface to the tangent plane on the spherical camera is given by (see Gårding and Lindeberg (1996) Equation (28), although here reformulated with somewhat different conventions and notation)
@@ 3.1.2 Binocular locally linearized projective projection
Figure 4(right) shows a schematic example of such a disparity field, as arising from a local linearization between corresponding points under the projective mapping between the two image domain for specific values of the geometric parameters $\mu$, $\gamma$ and $(Z_{X},Z_{Y})$.
@@
To characterize the effects that affine image transformations have on receptive field responses, let us next consider receptive fields according to the generalized Gaussian derivative model for visual receptive fields (Lindeberg 2021), and with this regard specifically only consider the effects in relation to the pure spatial and spatio-temporal smoothing transformations applied to static image data or video data, respectively.
@@ 4.1 Affine covariance for purely spatial receptive fields
provided that the spatial covariance matrices $\Sigma^{\prime}$ and $\Sigma$ in the two domains are related according to (see Equations (29) and (30) in Lindeberg and Gårding (1997))
@@
Given these raw purely smoothed components of the receptive field responses, corresponding receptive field responses in terms of spatial derivatives over the two domains, smoothed by affine Gaussian kernels with matching spatial covariance matrices $\Sigma$ and $\Sigma^{\prime}$, can, in turn, be related according to
@@
These relationships, thus, show that, provided that spatial covariance matrices are related in an appropriate way, according to an actual affine image transformation, it is then possible to match the receptive field responses perfectly between the two domains. For such matching to be possible in an actual situation, the actual value of the spatial covariance matrices, and hence the shapes of the corresponding affine Gaussian derivative based receptive fields, must thereby be adapted to the actual affine transformation. Specifically, if the input stimuli involve a variability over a certain subspace or subspace of the degrees of freedom of the affine transformations, then the family of receptive field shapes must also comprise a variability over a that corresponding subspace or subdomain of the parameter space of the affine Gaussian derivative based receptive fields, in order for it to be possible to perfectly match the outputs from the affine Gaussian derivative kernels between the two image domains.
@@ 4.2 Affine covariance for joint spatio-temporal receptive fields
denote the spatio-temporal smoothed scale-space representations in the two domains. Then, these joint spatio-temporal scale-space representations are related according to (see Lindeberg (2023) Equation (54))
@@ 5 Degrees of freedom in the affine Gaussian derivative model for visual receptive fields
    describes the overall spatial size of the receptive field,
@@
    describes the eccentricity or the degree of elongation of the receptive field, in the sense that deviations in this ratio from the rotationally symmetric special case when $\epsilon=1$ correspond to more elongated or anisotropic receptive fields, and
@@
    according to (19) in the affine image transformations,
@@
    of the affine transformation, which, in turn, represents the essential variability in the amount of non-uniform spatial stretching in the pure stretching transformation ${\cal D}$ according to (20) in the affine image transformations,
@@
Figure 8: Variability in the orientation of directional derivatives of affine Gaussian derivative receptive fields (for $\sigma_{1}=4$ and $\sigma_{2}=1$), with the orientation angle $\varphi$ increasing from left to right. (top row) First-order directional derivatives of affine Gaussian kernels. (bottom row) Second-order directional derivatives of affine Gaussian kernels. (Horizontal axes: image coordinate $x_{1}\in[-10,10]$. Vertical axes: image coordinate $x_{2}\in[-10,10]$.)
@@ 6.1 Restricted 3-D variability of regular affine Gaussian derivative kernels
Because the variability of the shapes of affine Gaussian receptive fields in three-dimensional, while the variability of the affine image transformations is four-dimensional, the variability of the shapes of the receptive fields according to the affine Gaussian derivative model used for formulating the explicit expressions for the idealized models for the receptive fields of simple cells according to Equations (43) and (44), with the directions for which the directional derivatives of the affine Gaussian kernels are computed in directions parallel with the eigendirections of the affine Gaussian kernels, cannot, in its original formulation, span the full variability of the affine group.
@@ 6.2 Possible extensions to a 4-D variability to enable matching of affine Gaussian derivative responses
An operational explanation why affine covariance cannot be achieved for the regular affine Gaussian derivative kernels, with the orientation for computing the directional derivative of that regular affine Gaussian derivative kernel being required to be aligned with a principal direction of the covariance matrix of the affine Gaussian kernel, is that if the affine image transformation comprises a non-uniform scaling transformation in a direction that not coincides with the principal directions of the spatial covariance matrix in the affine Gaussian smoothing kernel, then the transformed receptive field of a regular affine Gaussian derivative kernel, will be no longer within the definition of such a regular affine Gaussian directional derivative kernel.
@@
6.2.1 Alternative 1: Explicit expansion of the directions of directional derivatives of affine Gaussian kernels
The variability of the receptive fields of receptive fields according to the affine Gaussian derivative model for visual receptive fields can, however, be extended, by allowing the orientation angle for which the directional derivatives in the direction $\varphi$ of the affine Gaussian derivative kernels according to (42)
are computed, to instead be allowed to be computed in directions in the spatial domain that do not necessarily have to be aligned with either of the eigenvectors of the spatial covariance matrix $\Sigma$, see Figure 8 for examples of such receptive fields for fixed values of the spatial scale parameters $\sigma_{1}$ and $\sigma_{2}$. Then, such extended models for the spatial receptive fields, which will span a four-dimensional variability, will allow for expanded covariance properties over the full group of non-singular spatial affine transformations, according to the treatment in (Lindeberg 2024c Section 3.5).
@@
Thus, from a computational viewpoint, it is not necessary to fully span the variability of this fourth dimension, while nevertheless being able to match the receptive field responses between the two domains, if we would simultaneously expand the matching process to not just match plain receptive field responses between the two domains, but instead allow for matching of the receptive field response from the first image domain to a linear combination of receptive field responses over the second image domain of the form
@@
Given (i) these theoretical results, and given that (ii) the biological receptive fields corresponding to simple cells in the primary visual cortex, as measured by (1995; 2004), Conway and Livingstone (2006) and Johnson et al. (2008), can be qualitatively rather well modelled with spatial components in terms of affine Gaussian derivatives according to the axiomatically derived normative theory for visual receptive fields in Lindeberg (2013a; 2021), one may ask if biological vision has also developed corresponding variabilities in receptive field shapes as would be predicted from the presented theory. Specifically, one may ask if the shapes of the simple cells in the primary visual cortex would have the ability to span the variabilities corresponding to 2-D spatial affine transformations.

### v3

@@ 1 Introduction
For simplicity, we have here discarded a possible complementary variability with respect to an added translation vector $b$ in the image domain of the form $x^{\prime}={\cal A}\,x+b$, by without essential loss of generality assuming that all the multi-view observations of the same surface patch are focused on the same viewing point $P$ on the 3-D object. The origins of the coordinate systems in the two image domains are also assumed to correspond to the projections of that same fixation point.
@@
A conceptual question that one could then ask is if the ability of perceptual system to derive seemingly stable representations of external objects in the 3-D environment would lead to constraints regarding the variabilities in the shapes of the receptive fields at the earliest levels in the visual processing hierarchy. In a theoretically motivated and axiomatically determined normative theory of visual receptive fields (Lindeberg 2013a; 2021), idealized models for the receptive fields of simple cells in the primary visual cortex (V1) have specifically been derived. These receptive field models lead to provable covariance properties under spatial affine transformations, and regarding extensions to time-dependent spatio-temporal image data also provable covariance properties under Galilean transformations, to handle the variabilities in the relative motion between the objects in the world and the observer. See Lindeberg (2023; 2025b) for more in-depth treatments regarding the importance of covariance properties for visual receptive fields under geometric image transformations.
From a computational viewpoint, if we regard the earliest layers in the visual perception system as a computational engine, that infers successively more complex cues about the structure of the environment from the image measurements that reach the retina, then covariance properties of visual receptive fields do from a theoretical perspective constitute a highly useful notion. Specifically, if we design a computational algorithm for computing the local surface orientation of a smooth surface patch from binocular cues, then compared to not basing the computations on covariant spatial receptive fields, the use of provably affine-covariant visual receptive fields can improve the accuracy by an order of magnitude in estimates of the local surface orientation (see Tables 1–4 in Lindeberg and Gårding (1997)). This is achieved by eliminating a source of error corresponding to a mismatch between the backprojected receptive fields to the tangent plane of the surface, from the two different perspective views, see Figure 2 for an illustration.
@@
  • First theoretically describing the different degrees of freedom in spatial affine transformations, in terms of a canonical parameterization, based on a closed-form factorization of the affine transformation matrix ${\cal A}$ very closely related to a singular value decomposition.
  • Then, relating these degrees of freedom to the degrees of freedom in the parameters of the idealized receptive models of simple cells in the primary visual cortex, in terms of the generalized Gaussian derivative theory for visual receptive fields.
  • Finally, relating these results to existing evidence regarding variabilities in the shapes of biological receptive fields established from neurophysiological measurements.
Specifically, we will after an initial overview of related work in Section 2, in Section 3 describe a decomposition of 2-D spatial affine image transformations, based on a closed-form decomposition of the affine transformation matrix in a way very closely related to a singular value decomposition. The added value of this parameterization is that it reveals the degrees of freedom of 2-D spatial affine image transformations in a very geometric manner.
@@
Based on this theoretical basis, we will then in Section 6 relate the degrees of freedom in the affine Gaussian derivative model to the degrees of freedom revealed by our proposed decomposition of 2-D affine transformations. This is achieved in terms of explicit mappings between the parameters in our decomposition of 2-D spatial affine transformations and the parameters in the affine Gaussian derivative model. These connections will then be used in Section 7 for addressing the question of whether we can regard the receptive fields in the primary visual cortex as being able to span the degrees of freedom of 2-D affine transformations.
@@
A main significance of this study is that it provides a theoretical framework for how the variability in the spatial receptive field shapes in higher mammals can be accounted for based on symmetry properties of the environment. In particular, the theoretical relations to be derived reveal the practical usefulness for the visual system to explicitly handle geometric image transformations from the 3-D environment to the 2-D visual patterns registered in the earliest layers of the visual hierarchy.
More specifically, by expanding the receptive field shapes over the degrees of freedom of the geometric image transformations, the visual system will have the possibility to compute explicitly readable signals to the higher cortices, that are suitable for a variety of different imaging conditions, for extracting invariant features from the objects that are projected to the retina.
In this way, the visual system can meet the huge variability in spatial or spatio-temporal image structures, that are generated from different possible imaging conditions of any object in the environment by a corresponding variability in the receptive field shapes in the primary visual cortex. Thereby, it is maintaining a notion of identity between measurements of the same object performed under different geometric configurations between the object and the observer.
An important consequence of the covariance property of receptive fields, used as a theoretical foundation in following treatment, is that it makes it possible for receptive field responses to be matched between different viewing conditions. By the notion of geometric covariance, the visual system will be able to meet the variability in image structures generated by geometric image transformations by a corresponding variability in receptive field shapes. This implies that the receptive field responses will, to first order of approximation, be the same under the geometric image transformations, and thus achieving a constancy of the representation, given that the parameters of receptive fields can be matched to the geometric image transformations.
Thus, in summary, the proposed theory predicts a set of variabilities in receptive field shapes as induced by corresponding variabilities in image structures, and as generated by the variabilities of geometric image transformations.
@@ 2 Relations to previous work
The receptive field shapes of simple cells have, in turn, been modelled mathematically in terms of, most commonly, either Gabor filters by Marcelja (1980), Jones and Palmer (1987a; 1987b), Porat and Zeevi (1988), Ringach (2002; 2004), or Gaussian derivatives by Koenderink and van Doorn (1984; 1987; 1992) Young and his co-workers (1987; 2001; 2001) and Lindeberg (2013a; 2021). See specifically Figures 16 and 17 in Lindeberg (2021) for examples of how spatial receptive fields of simple cells in the primary visual cortex can be modelled in terms of affine Gaussian derivatives.
@@
The subject of this paper is to address this issue with regard to the special case of spatial affine image transformations, and specifically based on a canonical decomposition of such 2-D spatial affine transformations. This decomposition is in terms of a product form of (i) a uniform scaling transformation, (ii) pure rotations, (iii) a non-uniform scaling transformation, and (iv) a normalization of the non-uniform scaling transformation to a preferred symmetry orientation in the image domain.
An earlier study of parameterizations of affine transformations, which we will build upon in this work, was presented in (Lindeberg 1995). In this paper, we extend that parameterization by first of all providing an explicit derivation in Appendix A.1, which was not presented in the original paper because of space limitations. Then, we analyse the properties of this decomposition for four basic classes of primitive affine transformations in Appendix A.2. We do additionally give a more general treatment regarding special cases for the parameters in the affine transformation model, which thereby increases the domain of applicability of the proposed parameterization to more general families of affine transformation matrices, that are reasonably near the unit matrix multiplied by a scalar scaling factor.
@@ 3 The degrees of freedom in 2-D spatial affine transformations
and parameterize the image coordinates as $x=(x_{1},x_{2})^{T}$ and $x^{\prime}=(x^{\prime}_{1},x^{\prime}_{2})^{T}$. Here, we are specifically interested in the special case when the affine transformation matrix ${\cal A}$ is reasonably close to the unit matrix ${\cal I}$ multiplied by some scalar scaling factor.
@@ 3.1 Geometric interpretations of affine image deformations
As previously mentioned, let us consider surface patterns of possibly moving smooth objects in the world are projected to the image plane, by either (i) the perspective mapping to a monocular image, (ii) the projective mapping between two binocular views, or (iii) the temporally integrated optic flow fields obtained by observing moving objects between adjacent time moments. Then, the resulting image transformations or deformation fields can to first order of approximation be modelled as local affine transformations.
For this purpose, the decomposition in (LABEL:eq-aff-decomp) constitutes a geometrically very natural way to parameterize the degrees of freedom in such local affine transformations of the perspective/projective mappings. Specifically, these geometric image transformations determine the variabilities of the image structures, that a visual system is exposed to, when observing sufficiently smooth surface patterns in the environment from different viewing directions relative to the objects in the environment.
@@ 3.1.1 Monocular locally linearized perspective projection
To make such relations more explicit, consider, for example, the monocular perspective projection of a smooth local surface patch onto a spherical camera. Let us specifically consider two local coordinate systems, with
@@
  • the second local coordinate system being in the tangent plane of the spherical camera, with the second local coordinate direction aligned with the backprojected tilt direction onto the resulting image plane parallel to the tangent plane of the surface.
Then, the first-order locally linearized mapping from the tangent plane of the surface to the tangent plane on the spherical camera is given by (see Gårding and Lindeberg (1996) Equation (28), although here reformulated with somewhat different conventions and notation)
@@ 3.1.2 Binocular locally linearized projective projection
Figure 4(right) shows a schematic example of such a disparity field, as arising from a local linearization between corresponding points under the projective mapping between the two image domains (seen from the two views) for specific values of the geometric parameters $\mu$, $\gamma$ and $(Z_{X},Z_{Y})$.
@@
To characterize the effects that affine image transformations have on receptive field responses, let us next consider receptive fields according to the generalized Gaussian derivative model for visual receptive fields (Lindeberg 2021). With this regard, let us specifically only consider the effects in relation to the pure spatial and spatio-temporal smoothing transformations applied to static image data or video data, respectively.
@@ 4.1 Affine covariance for purely spatial receptive fields
provided that the spatial covariance matrices $\Sigma^{\prime}$ and $\Sigma$ in the two images domains (before and after the geometric image transformation) are related according to (see Equations (29) and (30) in Lindeberg and Gårding (1997))
@@
Given these raw purely smoothed components of the receptive field responses, corresponding receptive field responses in terms of spatial derivatives over the two image domains (before and after the geometric image transformation), smoothed by affine Gaussian kernels with matching spatial covariance matrices $\Sigma$ and $\Sigma^{\prime}$, can, in turn, be related according to
@@
These relationships, thus, show that, provided that spatial covariance matrices are related in an appropriate way, according to an actual affine image transformation, it is then possible to match the receptive field responses perfectly between the two image domains. For such matching to be possible in an actual situation, the actual value of the spatial covariance matrices, and hence the shapes of the corresponding affine Gaussian derivative based receptive fields, must thereby be adapted to the actual affine transformation. Specifically, if the input stimuli involve a variability over a certain subspace or subspace of the degrees of freedom of the affine transformations, then the family of receptive field shapes must also comprise a variability over a that corresponding subspace or subdomain of the parameter space of the affine Gaussian derivative based receptive fields, in order for it to be possible to perfectly match the outputs from the affine Gaussian derivative kernels between the two image domains.
@@ 4.2 Affine covariance for joint spatio-temporal receptive fields
denote the spatio-temporal smoothed scale-space representations in the two spatio-temporal image domains. Then, these joint spatio-temporal scale-space representations are related according to (see Lindeberg (2023) Equation (54))
@@ 5 Degrees of freedom in the affine Gaussian derivative model for visual receptive fields
    describes the overall spatial size of the receptive field.
@@
    describes the eccentricity or the degree of elongation of the receptive field, in the sense that deviations in this ratio from the rotationally symmetric special case when $\epsilon=1$ correspond to more elongated or anisotropic receptive fields.
@@
    according to (19) in the affine image transformations.
@@
    of the affine transformation, which, in turn, represents the essential variability in the amount of non-uniform spatial stretching in the pure stretching transformation ${\cal D}$ according to (20) in the affine image transformations.
@@
Figure 8: Variability in the orientation of directional derivatives of affine Gaussian derivative receptive fields (for $\sigma_{1}=4$ and $\sigma_{2}=1$ and preferred image orientation $\varphi=0$), with the orientation angle $\varphi$ for the directional derivative operators increasing from left to right. (top row) First-order directional derivatives of affine Gaussian kernels. (bottom row) Second-order directional derivatives of affine Gaussian kernels. (Horizontal axes: image coordinate $x_{1}\in[-10,10]$. Vertical axes: image coordinate $x_{2}\in[-10,10]$.)
@@ 6.1 Restricted 3-D variability of regular affine Gaussian derivative kernels
Consider the affine Gaussian derivative model used for formulating the explicit expressions for the idealized models for the receptive fields of simple cells according to Equations (43) and (44), with the directions for which the directional derivatives of the affine Gaussian kernels are computed in directions parallel with the eigendirections of the affine Gaussian kernels. Because the variability of the resulting shapes of affine Gaussian derivative-based receptive fields in three-dimensional, while the variability of the affine image transformations is four-dimensional, the variability of the shapes of the receptive fields according that regular affine Gaussian derivative model, in its original formulation, cannot span the full variability of the affine group.
@@ 6.2 Possible extensions to a 4-D variability to enable matching of affine Gaussian derivative responses
An operational explanation, why affine covariance cannot be achieved for the regular affine Gaussian derivative kernels, with the orientation for computing the directional derivative of that regular affine Gaussian derivative kernel being required to be aligned with a principal direction of the covariance matrix of the affine Gaussian kernel, can be stated as follows: If the affine image transformation comprises a non-uniform scaling transformation in a direction that not coincides with the principal directions of the spatial covariance matrix in the affine Gaussian smoothing kernel, then the transformed receptive field of a regular affine Gaussian derivative kernel, will be no longer within the definition of such a regular affine Gaussian directional derivative kernel.
@@
6.2.1 Alternative 1: Explicit expansion over the directions of directional derivatives of affine Gaussian kernels
The variability of the receptive fields of receptive fields according to the affine Gaussian derivative model for visual receptive fields can, however, be extended. Consider the orientation angle for which the directional derivatives in the direction $\varphi$ of the affine Gaussian derivative kernels are computed according to (42)
$T_{\scriptsize\mbox{simple}}(x_{1},x_{2};\;\sigma_{1},\sigma_{2},\varphi,m)=\sigma_{1}^{m}\,\partial_{\varphi}^{m}\,g(x_{1},x_{2};\;\Sigma).$ (51)
Let us instead allow the directional derivatives to be computed in directions $\varphi$ in the spatial domain that do not necessarily have to be aligned with either of the eigenvectors of the spatial covariance matrix $\Sigma$, see Figure 8 for examples of such receptive fields for fixed values of the spatial scale parameters $\sigma_{1}$ and $\sigma_{2}$. Then, such extended models for the spatial receptive fields, which will span a four-dimensional variability, will allow for expanded covariance properties over the full group of non-singular spatial affine transformations, according to the treatment in (Lindeberg 2025b Section 3.5).
@@
Thus, from a computational viewpoint, it is not necessary to fully span the variability of this fourth dimension, while nevertheless being able to match the receptive field responses between the two images domains, before and after the geometric image transformation. This can be achieved by simultaneously expanding the matching process to not just match plain receptive field responses between the two image domains, but instead allowing for matching of the receptive field response from the first image domain to a linear combination of receptive field responses over the second image domain of the form
@@
Given (i) these theoretical results, and given that (ii) the biological receptive fields corresponding to simple cells in the primary visual cortex, as measured by (1995; 2004), Conway and Livingstone (2006) and Johnson et al. (2008), can be qualitatively rather well modelled with spatial components in terms of affine Gaussian derivatives according to the axiomatically derived normative theory for visual receptive fields in Lindeberg (2013a; 2021), one may ask if biological vision has also developed corresponding variabilities in receptive field shapes as would be predicted from the presented theory? Specifically, one may ask if the shapes of the simple cells in the primary visual cortex would have the ability to span the variabilities corresponding to 2-D spatial affine transformations?